Two-point charges reacted with each other to produce 4.32 x 10⁻⁴ J. the magnitude of q is 30 nC with a distance from the point charge of 20 cm. Find the magnitude of the point charge



1
Expert's answer
2022-03-18T09:07:58-0400

The potential energy of interaction of the charges

"E_p=k\frac{q_1q_2}{r}"

Hence, the magnitude of the second charge

"q_2=\frac{E_p r}{kq_1}"

"q_2=\frac{4.32\times 10^{-4}\times 0.20}{9\times 10^9\times 30\times 10^{-9}}=3.2\times 10^{-7}\:\rm C"